OpenStudy (anonymous):

.5(160-x)= 50 how do i solve??

geerky42 (geerky42):

By have x be alone. \[\dfrac{1}{2}(160-x) = 50\] Seems we need to get rid of \(\dfrac{1}{2}\) first. Do you know how?

OpenStudy (anonymous):

no